# An Overview on Over-the-air Electromagnetic Signal Processing

### v2

@@ VI Discussion and Future Research Directions
In the previous sections, we explored the significant advancements in the development of ESP in recent years. By processing the signal directly at the EM level through the interaction of passive scatterers, ESP offers considerable benefits, including reduced power consumption, lower latency, and smaller device size compared to fully digital or hybrid alternatives. These features make ESP-based solutions highly promising for addressing scalability and sustainability challenges. However, these advantages are accompanied by trade-offs. New challenges must be considered, as they may limit the processing flexibility and applicability of these solutions compared to traditional digital-based implementations. Specifically, signal processing schemes must account for constraints imposed by fundamental EM laws, and the characteristics of the specific technology employed can further restrict the DoF available for implementing certain functionalities, as demonstrated in the previous sections. Therefore, it is crucial to adopt physically consistent models and design tailored EM-compliant signal processing methods.
@@
  • Optimization algorithms: From the SP perspective, the primary challenges lie in developing efficient techniques to address high-dimensional optimization problems constrained by EM laws, such as that expressed in (40), particularly when dealing with a large number of reconfigurable scattering elements. To achieve an improved balance between power consumption, scalability, latency, and processing flexibility, hybrid schemes that jointly optimize the digital and EM sections can also be explored. Another open challenge involves designing optimization techniques capable of determining the optimal geometric configuration of scatterers to achieve a desired class of signal processing functionalities.
  • Space-time-frequency EM processing: Most of the existing literature has focused on linear, time-invariant ESP solutions that operate primarily in the spatial domain. Consequently, our analysis in the previous sections has been largely confined to narrowband and time-invariant systems. However, emerging research is exploring the extension of these concepts to wideband and time-varying systems, addressing processing across the space, time, and frequency domains. For instance, EM devices operating with wideband signals often exhibit frequency-dependent effects, such as beam squinting or unwanted phase variations. While traditionally viewed as limitations, these phenomena can be strategically harnessed to enhance EM-level functionalities. By deliberately exploiting frequency-dependent behaviors, it becomes possible to design systems capable of tailored responses for signals at different frequencies. This approach is particularly relevant in the study of space-time modulated (STM)-RIS, also known as space-time modulated metasurfaces (STMM), which integrate phase and temporal phase gradients across metasurfaces to enable anomalous, non-reciprocal reflections [59]. This working principle generalizes the design concepts of traditional RIS.
    Another promising research direction involves integrating these concepts with optical technologies, potentially combining the strengths of both domains to unlock new capabilities.
@@
  • Artificial intelligence (AI)-aided ESP and ESP-aided AI: To address the so-called curse of dimensionality in reconfigurable scatterer optimization, AI-aided schemes that incorporate physically consistent models can be employed. Additionally, AI-based methods can be leveraged to develop hybrid solutions, where AI-driven optimization techniques are combined with traditional model-based approaches. This integration can lead to the creation of more efficient and adaptive ESP systems, capable of handling complex and dynamic signal processing tasks. Conversely, non-linear elements can be incorporated into reconfigurable scattering devices, such as SIMs, to emulate the behavior of a DNN through the EM device. However, unlike conventional DNNs, where weights and activation functions can be freely chosen, EM physical and technological constraints impose additional limitations. Consequently, models and training schemes must be reimagined to accommodate these constraints.

### v3

@@ VI Discussion and Future Research Directions
In previous sections, we explored the significant advancements in the development of ESP in recent years. By processing the signal directly at the EM level through the interaction of passive scatterers, ESP offers considerable benefits, including reduced power consumption, lower latency, and smaller device size compared to fully digital or hybrid alternatives. These features make ESP-based solutions highly promising for addressing scalability and sustainability challenges. However, these advantages are accompanied by trade-offs. New challenges must be considered, as they may limit the processing flexibility and applicability of these solutions compared to traditional digital-based implementations. Specifically, signal processing schemes must account for constraints imposed by fundamental EM laws, and the characteristics of the specific technology employed can further restrict the DoF available for implementing certain functionalities, as demonstrated in the previous sections. Therefore, it is crucial to adopt physically consistent models, such as those based on multi-port circuits, and design tailored EM-compliant signal processing methods.
@@
Figure 14: Left; The tri-hybrid architecture for multi-domain processing. Right: Example of a nonlinear SIM.
  • Optimization algorithms and novel hybrid SP architectures: From the SP perspective, an important issue is the development of efficient techniques to address high-dimensional optimization problems constrained by EM laws, such as that expressed in (42), particularly when dealing with a large number of reconfigurable scattering elements. Evidently, ESP alone cannot address all issues. A promising approach is the so-called tri-hybrid architecture [8] shown in Fig. 14-left, where SP is distributed across digital, analog (e.g., through phase shifters and adders), and EM domains. From the SP algorithm design perspective, the key task is to determine how to partition processing among these domains to achieve a given trade-off between power consumption, scalability, latency, and flexibility. Another open challenge involves designing optimization techniques capable of determining the optimal geometric configuration of scatterers to achieve a desired class of signal processing functionalities.
  • Space-time-frequency EM processing: Most of the existing literature has focused on linear, time-invariant ESP solutions that operate primarily in the spatial domain. Consequently, our analysis in the previous sections has been largely confined to narrowband and time-invariant systems. However, emerging research is exploring the extension of these concepts to wideband and time-varying systems, addressing processing across the space, time, and frequency domains. For instance, EM devices operating with wideband signals often exhibit frequency-dependent effects, such as beam squinting or unwanted phase variations. While traditionally viewed as limitations, these phenomena can be strategically harnessed to enhance EM-level functionalities. By deliberately exploiting frequency-dependent behaviors, it becomes possible to design systems capable of tailored responses for signals at different frequencies. This approach is particularly relevant in the study of space-time modulated (STM)-RIS, also known as space-time modulated metasurfaces (STMM), which integrates phase and temporal phase gradients across metasurfaces to enable anomalous, non-reciprocal reflections [58]. This working principle generalizes the design concepts of traditional RIS.
@@
  • Artificial intelligence (AI)-aided ESP and ESP-aided AI: To address the so-called curse of dimensionality in reconfigurable scatterer optimization, AI-aided schemes that incorporate physically consistent models can be employed. Additionally, AI-based methods can be leveraged to develop hybrid solutions, where AI-driven optimization techniques are combined with traditional model-based approaches. This integration can lead to the creation of more efficient and adaptive ESP systems, capable of handling complex and dynamic signal processing tasks. Conversely, passive non-linear elements can be incorporated into reconfigurable scattering devices, such as SIMs, to emulate the behavior of a DNN through the EM device. For instance, in Fig. 14-right, a layer containing meta-atoms with nonlinear scattering properties is inserted between the linear layers of the SIM [60]. However, unlike conventional DNNs, where weights and activation functions can be freely chosen, EM physical and technological constraints impose additional limitations. Consequently, models and training schemes must be reimagined to accommodate these constraints. Morover, the fundamental question of under which conditions the EM device can act as a universal approximator has still to be answered.